# Dynamic Neural Curiosity Enhances Learning Flexibility for Autonomous Goal Discovery

### v1

@@ V-D Tonic and phasic activation
The slow boost component is the core of the tonic and phasic activation within the Locus Coeruleus model. In Figure 15, we monitored the activations within the tonic and phasic NF. During exploration, the tonic activation remains sustained as long as the robot is not habituated to the object. Once the learning begins, the slow boost component periodically reset and delivers excitation to the tonic and phasic NF. The rate at which the boost reset depends on the persistence mechanism. In accordance with previous research [55] who suggests that the error prediction contributes to the phasic activation within the LC, our model propose an indirect contribution of the error through the persistence mechanism. Indeed, the persistence architecture modulates these periods depending on the goal error (see Appendix -D, Figures 19 and 20 - bottom row). If the error of a goal is significant, the period between two selections is lasting longer. At the first goal selection for the baseline experiment, the boost delivers excitation for 16 seconds before dropping down. At the end of learning the same skill, the boost provides excitation for 6.5 seconds. The transient action MT node also influences the period of boost activation. The $\tau_{+}$ parameter and the frequency of selecting a goal are inversely related. If $\tau_{+}$ is large, then the goal selection happens less often. In our model, the phasic activation within LC is the consequence of the goal selection for learning.
@@ V-E A loop for incremental learning
Figure 16: Time course of object discovery and learning by introducing the cube first, the cylinder in second, and the ball last. Forward model error and learning progress depict the learning.
The architecture detects bottom-up events such as the presence of new objects in the scene and begins to explore through goal discovery. Here, we reduce the habituation time ($\tau_{+}$ in visual memory trace) so each object has a short exploration. We begin by introducing the cube, let the robot discover and learn a goal, then reproduce this step for the cylinder and the ball. This protocol exhibits the properties of the attentional system to focus on new stimuli. The time course of the experiment with the learning progress and forward model error is detailed in Figure 16. It is also possible to drive the robot into exploration again after habituation toward a specific location by increasing the object stimuli. In practice, it is sufficient to introduce a supplementary excitation within the object selection NF (habituation mechanism).
@@ VI Conclusion and Discussion
The autonomous discovery and learning of goals in robotics is a complex issue to address and the state of the art represents a wide spectrum of approaches. Here, we take inspiration from neuroscience by rooting curiosity with attention in order to allow a robotic arm to discover and learn different skills. More precisely, we propose a simple model of the Locus Coeruleus, a nucleus in the brain controlling the switch between exploration and task engagement via the influence of curiosity, habituation and persistence. At first, the tonic activation within the LC drives the robot to generate actions that will produce new stimuli, and thus discover new goals via bottom-up attention. Then, the curiosity mechanism directly influences the Locus Coeruleus activation with a phasic mode by engaging in learning, depending on both forward model error and learning progress of these goals. In addition, we introduce new cognitive mechanisms to regulate the discovery and learning of goals with respectively a habituation and persistence component. The generation of robot motion is performed by inhibition of return and dynamic movement primitives. The complete architecture is modelled with dynamic neural fields (DNFs) except for the forward and inverse models, which are multi-layer perceptrons. The use of DNFs supports several technical novelties such as the generation of poses around an object and the interaction between skills in memory. In addition, the goal error modelling allows widespread connections that bootstrap the shift between goal discovery and learning as well as switch the learning of goals. Regarding the habituation paradigm driving exploration, we demonstrated that the robot was able to discover various goals, with a higher average number of goals for a slow habituation. For persistence, we tuned the time spent on learning a goal as well as the strength of the error inhibition to the LC and demonstrated a clear influence on learning. Varying these parameters manifests positive and negative impacts on learning that suggest an individual optimization depending on the objects. Moreover, the curiosity module exhibits an interesting property in case of two similar goals where the learning of one directly inhibits the other. Finally, the results indicate that the architecture continuously oscillates between goal discovery through bottom-up attention and goal learning via curiosity.
@@
Nevertheless, the current model suffers from several limitations. The objects in habituation are distinguished base on colors, which is not realistic. In the future, we plan to use a variational autoencoder (VAE) to integrate more features and improve exploration. For the former, an object and a goal will be characterized by more features (3D object pose, touch sensor, rotation), so a more diverse set of skills can be learned. Regarding the latter, we will use the latent space as a supplementary method for discovering new goals through direct exploration [37]. This way, we will investigate the balance between direct exploration (VAE) and random exploration with the inhibition of return method. The end-effector orientation as well as the $z$ dimension of the robot were fixed and we intend to remove this limitation to perform a diverse range of robot actions that could lead to different outcomes and produce a developmental trajectory. Then, we will explore the tuning of the persistence and error inhibition parameters together. Indeed, there are several possibilities offered by the model and different learning patterns could still be observed. We would like to investigate if the number of goals discovered during habituation can be used to predict the parameters for an optimal learning. Finally, The goals discovered were on the same developmental level and the first learning selection entirely depended on the weights initialization of the forward models. We intend to investigate if the persistence mechanism can autonomously shift from a short to a long persistence in order to bootstrap the learning progress of the discovered goals.

### v2

@@ V-D Tonic and phasic activation
The slow boost component is the core of tonic and phasic activation within the Locus Coeruleus model. In Figure 9, we monitored activations within the tonic and phasic NF. During exploration, the tonic activation remains sustained as long as the robot is not habituated to the object. After learning begins, the slow boost component is periodically reset and provides excitation to the tonic and phasic NF. The boost reset rate depends on the persistence mechanism. Our model proposes an indirect contribution of the error through the persistence mechanism. Indeed, the persistence architecture modulates these periods depending on the goal error (see Appendix -D, Figures 13 and 14 - bottom row). When the error of a goal is substantial, the interval between successive selections increases. In the first goal selection for the baseline experiment, the boost delivers excitation for 16 seconds before dropping down. At the end of learning the same goal, the boost provides excitation for 6.5 seconds. The transient action MT node also influences the period of boost activation. The parameter $\tau_{+}$ and the frequency of selecting a goal are inversely related. If $\tau_{+}$ is large, then the selection of goals occurs less often. Moreover, error inhibition leads to more frequent disengagement if there is no learning progress. In our model, the phasic activation of LC results from the goal selection process and the frequency with which the robot switches the learning between different goals.
@@ V-E A loop for incremental learning
Figure 10: Timecourse of object discovery and learning by introducing the cube first, the cylinder in second, and the ball last. Forward model error and learning progress depict the learning trajectory.
The architecture detects bottom-up events, such as the presence of new objects in the scene, and begins to explore through goal discovery. Here, we reduce the habituation time ($\tau_{+}$ in visual memory trace), so that each object has a short exploration. We begin by introducing the cube, let the robot discover and learn a goal, then reproduce this step for the cylinder and the ball. This protocol exhibits the properties of the attentional system to focus on new stimuli. The time course of the experiment with the learning progress and forward model error is detailed in Figure 10. It is also possible to drive the robot into exploration again after habituation to a specific location by increasing the object stimuli. In practice, it is sufficient to introduce a supplementary excitation within the object selection NF in the habituation mechanism.
@@ VI Conclusion and Discussion
The autonomous discovery and learning of goals in robotics is a complex issue to address, and the state-of-the-art represents a wide spectrum of approaches. Here, we take inspiration from neuroscience by rooting curiosity with attention in order to allow a robotic arm to discover and learn different goals. More precisely, we propose a simple model of the Locus Coeruleus, a nucleus in the brain that controls to some extent the switch between exploration and task engagement through the influence of curiosity, habituation, and persistence. At first, the tonic activation within the LC drives the robot to generate actions that will produce new stimuli, and thus discover new goals via bottom-up attention. Then, the curiosity mechanism directly influences the Locus Coeruleus activation with a phasic mode by engaging in learning, depending on both forward model error and learning progress of these goals. In addition, we introduce new cognitive mechanisms to regulate the discovery and learning of goals with, respectively, a habituation and persistence component. The generation of robot motion is performed by inhibition of return and dynamic movement primitives. The complete architecture is modeled with dynamic neural fields (DNFs) except for the forward and inverse models, which are multilayer perceptrons. The use of DNFs supports several technical novelties, such as the generation of poses around an object and the interaction between goals in memory. In addition, the goal error modeling allows widespread connections that bootstrap the shift between goal discovery and learning, as well as control the learning of goals. Regarding the habituation paradigm driving exploration, we demonstrated that the robot was able to discover various goals, with a higher average number of goals for slow habituation. For persistence, we tuned the time spent learning a goal as well as the strength of the error inhibition to the LC and demonstrated a clear influence on learning. Varying these parameters manifests positive and negative impacts on learning that suggest individual optimization depending on the objects. Moreover, the curiosity module exhibits an interesting property in the case of two similar goals, where the learning of one directly inhibits the other. Finally, the results indicate that the architecture continuously oscillates between goal discovery through bottom-up attention and goal learning via curiosity.
@@
However, the current model suffers from several limitations. The objects in the habituation mechanism are distinguished by color, which is not realistic. In the future, we plan to use a variational autoencoder (VAE) to integrate more features and improve the discovery of goals, as demonstrated by researchers [84]. To do so, an object and a goal will be characterized by more features (3D object pose, touch sensor, rotation), so a more diverse set of goals can be learned. Regarding exploration, we will use the latent space as a supplementary method to discover new goals through direct exploration [42]. This way, we will investigate the balance between direct exploration (VAE) and random exploration with the inhibition-of-return method. The orientation of the end effector and the $z$ dimension of the robot were fixed, and we intend to eliminate this limitation to perform a diverse range of robot actions that could lead to different outcomes and potentially produce a developmental trajectory.
Finally, we will explore the tuning of the persistence and error-inhibition parameters together. In fact, the model offers several possibilities, and different learning patterns could still be observed. We would like to investigate whether the number of goals discovered during habituation can be used to predict the parameters for optimal learning, and how to autonomously determine and set these parameters. Finally, the goals discovered were on the same developmental level, and the first learning selection entirely depended on the weights initialization of the forward models. We intend to investigate the possibility of autonomously tuning the persistence and error-inhibition parameters to produce optimal learning based on the object difficulty.